OpenStudy (anonymous):

when n is divided by 12,the remainder is 6.What is the remainder when n is divided by 6?

OpenStudy (anonymous):

12/2= 6 6/2=3

OpenStudy (anonymous):

n is 2

OpenStudy (anonymous):

i didnt get u

OpenStudy (anonymous):

how u solve it can u do it step by step @dread56ace

OpenStudy (anonymous):

yes plz @dread56ace

OpenStudy (anonymous):

when u divide n by 12 you get a remainder of 6 right. so what number do you divide from 12 to get 6

OpenStudy (anonymous):

6

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@Henry.Lister do u know how to solve it?

OpenStudy (anonymous):

what ur not understanding you need to find what n is

OpenStudy (anonymous):

n/12 = 6 n = 72 therefore; 72/6 = 12

OpenStudy (amriju):

take an integer K...then we can write n as n=12*K+6....K can be changed depending on n...so now we divide n by 6 thats (12*K+6)/6=2K+1...remainder is 0

OpenStudy (amriju):

its remainder not quotient @pathosdebater

OpenStudy (anonymous):

If the quotient is a whole number, then the remainder is obviously 0... I was showing how to find that out

OpenStudy (amriju):

note that 12K+6...always leaves a remainder 6 when divided by 12..this is the genaral form of the integer n

OpenStudy (amriju):

hey...but n/12 not=6 as u hav written ....u must hav overlookd @pathosdebater

OpenStudy (anonymous):

what if the quotient is not a whole no.?

OpenStudy (amriju):

then u have a remainder not equal to 0

OpenStudy (anonymous):

So n = 78

OpenStudy (amriju):

n cant have any specific value...what about n=114...divide it by 12....u hav a remainder 6...@pathosdebater

OpenStudy (anonymous):

True

OpenStudy (amriju):

n=12*k+6..as i mentioned...k is any integer...so put k=1,2,3.....and u 'll get diff values of n

OpenStudy (anonymous):

what about this one @amriju when n is divided by 5,the remainder is 6.What is the remainder when n is divided by 7?

OpenStudy (amriju):

did u make up the ques..?...the ques is wrong..

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i just want to know that if the quotient is not a whole no.then what wud be the solution of that question?

OpenStudy (amriju):

look...if the quetiont is not a whole number.. then u have a remainder...thats all...example..114/12...12*9=108..u have remainder 6..so the quo. is..9+(6/12)...thats 9.5 not awhole num...but 8/4....we have 4*2=8..rem=0...so quo=2+(0/4)=2..whole number

OpenStudy (anonymous):

Examples: 12/5 leaves a remainder of 2. (12+4)/5 leaves a remainder of 1.

OpenStudy (anonymous):

(x^3 - 2x^2 + 4x - 2)/(x-3) leaves a remainder of 19.

OpenStudy (anonymous):

i got it thanku